4 An elastic collision occurs in one dimension, in which a 10 kg block traveling at 5 m/s collides with a 5 kg block traveling at 3 m/s in the same direction. What are the velocities of the two blocks immediately after the collision?

4

An elastic collision occurs in one dimension, in which a 10 kg block traveling at 5 m/s collides with a 5 kg block traveling at 3 m/s in the same direction. What are the velocities of the two blocks immediately after the collision?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
let's consider velocities of blocks after collision is V1' and V2'
m1 = 10 kg, m2 = 5 kg. V1 = 5 m/s, V2 = 3 m/s
from the conservation of momentum
mV1+m2V2 = m1V1'+m2V2'
10*5+5*3 = 10*V1'+5*V2'
65 = 10V1'+5V2'----------------------------(i)
from the conservation of energy
1/2*10*(5)^2+1/2*5*(3)^2 = 1/2*10*V1'^2+1/2*5*V2'^2
250+45 = 10V1"2+5V2'^2
295 = 10V1'^2+5V2'^2--------------------------(ii)
On solving (i) and (ii)
v1' = - 1 m/s and V2' = 15 m/s
Thanks & Regards,
Nirmal Singh
Askiitians faculty